OpenStudy (amtran_bus):

Given AX^2+BX+c=0 Derive the general solution for x

OpenStudy (amtran_bus):

Would you use the quadratic formula?

OpenStudy (anonymous):

no, i don't think so

OpenStudy (anonymous):

have you learnt differentiation?

OpenStudy (amtran_bus):

To a degree, yes. This is actually a physics problem.

OpenStudy (amtran_bus):

Oh ok. The first D would be 2x or something, right?

OpenStudy (anonymous):

just wait... I've done a question like this last year for math methods, just trying to remember what I did

OpenStudy (amtran_bus):

your fine, thanks for trying to help.

OpenStudy (anonymous):

I know what the answer is, just trying to think of how to get there

OpenStudy (anonymous):

what you can try to do, instead of differentiation, is to complete the square. so, y=ax^2+bx+c y=a(x^2+bx/a+c/a) y=a(x^2+bx/a+(b/2a)^2-(b/2a)^2+c/a) y=a(x+b/2a)^2-a(b^2/4a^2+c/a) y=a(x+b/2a)^2+(c-b^2/4a)

OpenStudy (anonymous):

Confusing right?

OpenStudy (amtran_bus):

I think differentiation would be easier...

OpenStudy (anonymous):

Then using the form: y=a(x-b)+c turning point: (b,c) then: (-b/2a, c-b^2/4a)

OpenStudy (anonymous):

now let me try to figure it out differentiation style

OpenStudy (anonymous):

right!!! I sorta remember now!!!

OpenStudy (amtran_bus):

remembering is a good thing!

OpenStudy (anonymous):

1. you need to differentiate the equation y=ax^2+bx+c so, dy/dx=2ax+b

OpenStudy (amtran_bus):

Understand that step.

OpenStudy (anonymous):

then, you need to find the minimum of the graph or the turning point

OpenStudy (anonymous):

to do that, the 'gradient' or dy/dx must be zero

OpenStudy (anonymous):

so, dy/dx=2ax+b 0=2ax+b x=-b/2a

OpenStudy (anonymous):

since it is the general solution, it means the turning point...forgot to point that out

OpenStudy (amtran_bus):

Is that the critical point? If so, I understand, just beeb awhile.

OpenStudy (amtran_bus):

*been

OpenStudy (anonymous):

yes, the peak

OpenStudy (anonymous):

your answer would be x=-b/2a if you need to find y, just substitute -b/2a into the original equation y=ax^2+bx+c then you should get y=c-b^2/2a

OpenStudy (amtran_bus):

Is that it? Praise God! Thank you soooo much!

OpenStudy (anonymous):

sorry for the confusion earlier on, had to open my yr 11 textbook from last year!!!

OpenStudy (amtran_bus):

Oh my goodness. I'm a 12th grader in Tennessee...

OpenStudy (anonymous):

whoops!!! y=c-b^2/4a not y=c-b^2/2a

OpenStudy (anonymous):

no, I'm only an 11th grader :) I was doing advanced maths last year I'm from Australia

OpenStudy (amtran_bus):

Our American system needs help. Thanks!

OpenStudy (anonymous):

no, thank YOU!!! I needed the refreshment in differentiation

OpenStudy (anonymous):

Mind if I ask a few questions? I really don't mean to interrupt or cause confusion, but I wonder what you guys are trying to prove here, the general solutions for formula for a quadratic equation as given in the problem set? Or the equation of the min/max point.

OpenStudy (anonymous):

As it was said, the extrema is: \[\Large x=-\frac{b}{2a} \] Plugging that back into the original equation will give you the y-value of the extrema. Not the roots.

OpenStudy (amtran_bus):

Right, that is what I needed.

OpenStudy (amtran_bus):

That is, what chelsea04 provided.

OpenStudy (anonymous):

You need the roots? Or the extrema? Please, pardon me. But \[\Large ax^2+bx+c=0 \] is a quadratic equation, if it asks you to find the x, I read this question as: find the x-value that makes the given statement true. And not: Find the Maximum/Minimum / Zero-Slope.

OpenStudy (anonymous):

I am perfectly fine with @Chelsea04 proof, I just wonder if everyone was talking about the same thing here (-: I clearly misunderstood this problem then. Pardon me.

OpenStudy (amtran_bus):

Its cool! Have a great day :)

OpenStudy (amtran_bus):

@Spacelimbus can you tell me what the determinant is? I have a problem and need to find the determinant of a matrix is by hand.

OpenStudy (anonymous):

\[\Large \det(A)=ad-cb \] Considering the entries. \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]

OpenStudy (amtran_bus):

ohhh yea. Thanks! I should have made another question so I could have given you a best response!

OpenStudy (anonymous):

no worries at all, I am still wondering about the problem above *laughs* it just seems wrong to me to confuse the roots with the max.